# Nonlinear Schrödinger equation in terms of elliptic and hyperelliptic $\sigma$ functions

### 10 Mar 2024 v1

@@ Abstract.
It is known that the elliptic function solutions of the nonlinear Schrödinger equation are reduced to the algebraic differential relation in terms of the Weierstrass sigma function, ${\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha\frac{\partial}{\partial u}\right]\Psi+\frac{1}{2}\frac{\partial^{2}}{\partial u^{2}}\Psi+(\Psi^{*}\Psi)\Psi=\frac{1}{2}(2\beta+\wp(v)+\lambda_{2}-\alpha^{2})\Psi}$, where $\Psi(u;v,t):={\mathrm{e}}^{\alpha u+{\mathfrak{i}}\beta t+c}$ ${\frac{{\mathrm{e}}^{-\zeta(v)u}\sigma(u+v)}{\sigma(u)\sigma(v)}}$, its dual $\Psi^{*}(u;v,t)$, and certain complex numbers $\alpha,\beta$. In this paper, we generalize the algebraic differential relation to those of genera two and three in terms of the hyperelliptic sigma functions.
@@ 1. Introduction
(1.1) $\left[-{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha\frac{\partial}{\partial u}\right]\Psi+\frac{1}{2}\frac{\partial^{2}}{\partial u^{2}}\Psi+(\Psi^{*}\Psi)\Psi=\frac{1}{2}(2\beta+\wp(v)+\lambda_{2}-\alpha^{2})\Psi.$
@@
for the elliptic curve $E:y^{2}=(x-e_{1})(x-e_{2})(x-e_{3})$, $c\in{\mathbb{C}}$, and the Weierstrass sigma functions of $E$. After we fix the point $v$ and $e_{i}$ so that $\Psi^{*}$ is the complex conjugate of $\Psi$ and $(\alpha,u)$ belongs to ${\mathfrak{i}}{\mathbb{R}}\times{\mathbb{R}}$ or ×${\mathbb{R}}\times{\mathfrak{i}}{\mathbb{R}}$, we find the elliptic solutions of the NLS equation. We call these conditions the reality conditions of the NLS equation. Determination of the conditions depends on the situations which we require, and is not simple because we tune the moduli parameter $e_{i}$ as in [Ka].
@@
Construction of the algebraic solutions of the NLS equation was found by Previato 1985 [P] and is written in the book by Belokolos, Bobenko, Enolskii, Its and Matveev [BBEIM] precisely. On the other hand reevaluation of the theory of Kleinian sigma function, which was discovered by Klein have been performed by several authors [BEL1, BEL2, EEL, Ma6]. We have found the hyperelliptic solutions of the KdV equation [BEL1, Ma1], the MKdV equation [Ma1], the Toda lattice equation [KMP], the sine-Gordon equation [Ma3] in terms of the sigma functions and meromorphic functions of the symmetric product of algebraic curves. However the hyperelliptic solution of the NLS equation in terms of the Kleinian sigma functions has not been obtained yet. If we obtained it, it is easily to draw the numerical evaluations of the time development for a certain class of the curves [Ma5]. The approach is much easier than the approach by means of the the theta functions.
@@
In this paper, besides the reality conditions, we show the generalization of the algebraic differential relation (1.1) of genera two and three in terms of the hyperelliptic sigma functions associated with the hyperelliptic solutions of the NLS equation. To obtain this, we express the Baker function in terms of the sigma function reported in [BEL1], though our Baker function slightly differs from that in [BEL1]. Using the Baker equation, we show the generalization. We note that the Baker equation, or the generalized Lamé equation, of the Baker function contains the potential $\wp_{gg}$, while the addition theorem of the sigma functions gives the meromorphic functions of the Jacobian $J_{X}$, which does not coincide with $\wp_{gg}$ itself except for $g=1$. Thus, the discrepancy prevents us from obtaining the generalization of (1.1). Even for genera two and three, we find the generalization of the algebraic differential relation 1.1) in this paper. In other words, the purpose of this paper shows how to overcome the discrepancy so that we obtain the generalization of (1.1) associated with the hyperelliptic solution of the NLS equation.
@@
We note that in the NLS equation, we have crucial problems related to the reality conditions depending on the ”nonfocusing” and ”focusing” NLS equations. They correspond to the sign of the nonlinear terms in the NLS equation [P] and to the choice of $(\alpha,u)$ which belongs to ${\mathfrak{i}}{\mathbb{R}}\times{\mathbb{R}}$ or ×${\mathbb{R}}\times{\mathfrak{i}}{\mathbb{R}}$. However, we will not address these problems at this stage, as we plan to consider them in the next step.
@@
The contents of this paper are as follows. Section 2 is devoted for the genus one case or the derivation of (1.1). Section 3 shows the preliminary of the hyperelliptic sigma function. Section 4 provides the addition formula related to this NLS equation. In Section 5, we show the Baker functions in terms of the hyperelliptic sigma functions. In Section 6, we generalize (1.1) for the genera two and three cases as by using the Baker function and the addition formulae in Theorems 6.2 and 6.4.
@@ 2. Genus one case
$E:\ \frac{1}{4}{\widehat{y}}^{2}=y^{2}=x^{3}+\lambda_{2}x^{2}+\lambda_{1}x+\lambda_{0}$
[…]
where the $e$’s are complex numbers, $\lambda_{2}=-(e_{1}+e_{2}+e_{3})=0$ and ${\widehat{y}}=2y$.
@@
with $x(u)=\wp(u),~{}{\widehat{y}}(u)=\wp^{\prime}(u)$ and $\infty$ is the infinity point of $E$. The Jacobian is given by $J_{E}={\mathbb{C}}/(2{\mathbb{Z}}\omega^{\prime}+2{\mathbb{Z}}\omega^{\prime\prime})$ using the double period $(2\omega^{\prime},2\omega^{\prime\prime})$.
@@
Let $(\wp(u),\wp^{\prime}(u))=(x_{1},2y_{1})$ and $(\wp(v),\wp^{\prime}(v))=(x_{2},2y_{2})$. By using ${\frac{d}{du}=2y_{1}\frac{d}{dx_{1}}}$ and ${\frac{d}{dv}=2y_{2}\frac{d}{dx_{2}}}$, we find that (2.4) leads the following relations:
@@
We also have ${\displaystyle\frac{d\psi(u;v)}{du}}=(\zeta(u+v)-\zeta(u)-\zeta(v))\psi(u;v)$, or
@@
Further, by the Miura relation, $2x+\lambda_{2}+{\widetilde{x}}=\mu^{2}+\mu_{u}$ for ${\mu:=\frac{d}{du}\log\psi(u;v)=\frac{y-{\widetilde{y}}}{x-{\widetilde{x}}}}$, we have the static MKdV equation,
@@
$\Psi(u;v,t):={\mathrm{e}}^{\alpha u+{\mathfrak{i}}\beta t+c}\psi(u;v),\quad\Psi^{*}(u):=-{\mathrm{e}}^{-\alpha u-{\mathfrak{i}}\beta t-c}\psi(u;-v),$
for $\alpha,\beta,c\in{\mathbb{C}}$. Then the addition theorem (2.4) leads the relation,
(2.14) $\left[{\mathfrak{i}}\frac{\partial}{\partial t}-\alpha\frac{\partial}{\partial u}\right]\Psi+\frac{1}{2}\frac{\partial^{2}}{\partial u^{2}}\Psi+(\Psi^{*}\Psi)\Psi=\frac{1}{2}(2\beta-\wp(v)+\lambda_{2}-\alpha^{2})\Psi,$
which is the same as (1.1).
@@ 3. Hyperelliptic curve $X$ and sigma functions
In this Section, we give a background information on the hyperelliptic $\theta$ functions and the $\sigma$ functions as a generalization of the Weierstrass elliptic $\sigma$ function.
@@ 3.1. Geometrical setting of hyperelliptic curves
together with a smooth point $\infty$ at infinity. Here $\lambda$’s are complex numbers. The affine ring related to $X$ is denoted by $R_{X}:={\mathbb{C}}[x,y]/(y^{2}-f(x))$. We fix a basis of holomorphic one-forms
@@
$\omega^{\prime}=\frac{1}{2}\left[\oint_{\alpha_{j}}{\nu^{\mathrm{I}}_{i}}\right],\quad\omega^{\prime\prime}=\frac{1}{2}\left[\oint_{\beta_{j}}{\nu^{\mathrm{I}}_{i}}\right],$
@@
${\widetilde{w}}:S^{k}{\widetilde{X}}\to{\mathbb{C}},\quad{\widetilde{w}}(\gamma_{1},\ldots,\gamma_{k})=\sum_{i=1}^{k}\int_{\gamma_{i}}\begin{pmatrix}{\nu^{\mathrm{I}}_{1}}\\ \vdots\\ {\nu^{\mathrm{I}}_{g}}\end{pmatrix},$
$w:S^{k}X\to J_{X},\quad w((x_{1},y_{1}),\cdots,(x_{k},y_{k}))=\sum_{i=1}^{k}\int_{\infty}^{(x_{i},y_{i})}\begin{pmatrix}{\nu^{\mathrm{I}}_{1}}\\ \vdots\\ {\nu^{\mathrm{I}}_{g}}\end{pmatrix}\hbox{\rm mod}\ {\Lambda}.$
The image of $w$ is denoted by $W_{X}^{k}=w(S^{k}X)$, i.e., $J_{X}=W_{X}^{g}$. Further we introduce an injection $\iota_{X}:X\to{\widetilde{X}}$ and will fix it. We find that $w=\kappa_{J}\circ{\widetilde{w}}\circ\iota_{X}$.
@@
    ${\mathcal{X}}_{g}:={\small{\begin{pmatrix}\chi_{1,0}&\chi_{1,1}&\cdots&\chi_{1,g-1}\\ \chi_{2,0}&\chi_{2,1}&\cdots&\chi_{2,g-1}\\ \vdots&\vdots&\ddots&\vdots\\ \chi_{g,0}&\chi_{g,1}&\cdots&\chi_{g,g-1}\end{pmatrix}}},\quad{\mathcal{Y}}_{g}:={\small{\begin{pmatrix}y_{1}&\ &\ &\ \\ \ &y_{2}&\ &\ \\ \ &\ &\ddots&\ \\ \ &\ &\ &y_{g}\end{pmatrix}}}=\frac{1}{2}{\mathcal{H}}_{g},$
    ${\mathcal{F}}_{g}^{\prime}:={\small{\begin{pmatrix}F^{\prime}(x_{1})&&&\\ &F^{\prime}(x_{2})&&\\ &&\ddots&\\ &&&F^{\prime}(x_{g})\end{pmatrix}}},{\mathcal{V}}_{g}:={\small{\begin{pmatrix}1&1&\cdots&1\\ x_{1}&x_{2}&\cdots&x_{g}\\ x_{1}^{2}&x_{2}^{2}&\cdots&x_{g}^{2}\\ \cdot&\cdot&&\cdot\\ x_{1}^{g-1}&x_{2}^{g-1}&\cdots&x_{g}^{g-1}\end{pmatrix}}},$
    ${\mathcal{E}}_{g}:={\small{\begin{pmatrix}1&\ &\ &\ &\ \\ \varepsilon_{g-1}&1&\ &\text{\Large 0}&\ \\ \varepsilon_{g-2}&\varepsilon_{g-1}&1&&\ \\ \vdots&\vdots&\vdots&\ddots&\\ \varepsilon_{1}&\varepsilon_{2}&\cdots&\varepsilon_{g-1}&1\\ \end{pmatrix}}},\quad{\mathcal{K}}_{g}:={\small{\begin{pmatrix}x_{1}^{g-1}&x_{1}^{g-2}&\cdots&1\\ x_{2}^{g-1}&x_{2}^{g-2}&\cdots&1\\ \vdots&\vdots&\ddots&\vdots\\ x_{g}^{g-1}&x_{g}^{g-2}&\cdots&1\end{pmatrix}}}.$
@@
    The inverse matrix of ${\mathcal{X}}_{g}$ is given by ${\mathcal{X}}_{g}^{-1}={\mathcal{F}}_{g}^{\prime-1}{\mathcal{V}}_{g}$.
@@
    ${\small{\begin{pmatrix}\partial_{u_{1}}\\ \partial_{u_{2}}\\ \vdots\\ \partial_{u_{g}}\end{pmatrix}}}=2{\mathcal{Y}}_{g}{\mathcal{F}}_{g}^{\prime-1}{\mathcal{X}}_{g}{\small{\begin{pmatrix}\partial_{x_{1}}\\ \partial_{x_{2}}\\ \vdots\\ \partial_{x_{g}}\end{pmatrix},}}\quad\frac{\partial x_{i}}{\partial u_{r}}=\frac{2y_{i}}{F^{\prime}(x_{i})}\chi_{i,r-1},\quad\frac{\partial y_{i}}{\partial u_{r}}=\frac{f^{\prime}(x_{i})}{F^{\prime}(x_{i})}\chi_{i,r-1},$

### 14 Mar 2024 v2

@@ Abstract.
It is known that the elliptic function solutions of the nonlinear Schrödinger equation are reduced to the algebraic differential relation in terms of the Weierstrass sigma function, ${\left[-\mathfrak{i}\frac{\partial}{\partial t}+\alpha\frac{\partial}{\partial u}\right]\Psi-\frac{1}{2}\frac{\partial^{2}}{\partial u^{2}}\Psi+(\Psi^{*}\Psi)\Psi=\frac{1}{2}(2\beta+\alpha^{2}-3\wp(v))\Psi}$, where $\Psi(u;v,t):=\mathrm{e}^{\alpha u+{\mathfrak{i}}\beta t+c}$ ${\frac{\mathrm{e}^{-\zeta(v)u}\sigma(u+v)}{\sigma(u)\sigma(v)}}$, its dual $\Psi^{*}(u;v,t)$, and certain complex numbers $\alpha,\beta$ and $c$. In this paper, we generalize the algebraic differential relation to those of genera two and three in terms of the hyperelliptic sigma functions.
@@ 1. Introduction
(1.1) $\left[-{\mathfrak{i}}\frac{\partial}{\partial t}+\alpha\frac{\partial}{\partial u}\right]\Psi-\frac{1}{2}\frac{\partial^{2}}{\partial u^{2}}\Psi+(\Psi^{*}\Psi)\Psi=\frac{1}{2}(2\beta+\alpha^{2}-3\wp(v))\Psi,$
@@
for the elliptic curve $E:{\widehat{y}}^{2}=4(x-e_{1})(x-e_{2})(x-e_{3})$, $\alpha,\beta,c\in{\mathbb{C}}$, and the Weierstrass sigma functions, $\sigma$, and zeta function, $\zeta$, of $E$. After we fix the point $v$ and $e_{i}$ so that $\Psi^{*}$ is the complex conjugate of $\Psi$ and $(\alpha,u)$ belongs to ${\mathfrak{i}}{\mathbb{R}}\times{\mathbb{R}}$ or ×${\mathbb{R}}\times{\mathfrak{i}}{\mathbb{R}}$, we find the elliptic solutions of the NLS equation. We refer these conditions as the reality conditions of the NLS equation. Determination of the conditions depends on the situations which we require, and is not simple because we tune the moduli parameter $e_{i}$ as in [Ka].
@@
The construction of the algebraic solutions of the NLS equation was found by Previato 1985 [P] and is written in the book by Belokolos, Bobenko, Enolskii, Its and Matveev [BBEIM] precisely. On the other hand reevaluation of the theory of the hyperelliptic sigma function, which was discovered by Klein [Kl] and studied by Baker [B], has been performed by several authors [BEL1, BEL2, EEL, Ma6]. We found the hyperelliptic solutions of the KdV equation [BEL1, Ma1], MKdV equation [Ma1], Toda lattice equation [KMP], sine-Gordon equation [Ma3] in terms of the sigma functions and meromorphic functions of the symmetric product of algebraic curves. However, hyperelliptic solutions of the NLS equation in terms of the hyperelliptic sigma functions have not yet been obtained. When we obtain them, we can find the algebraic relations among the invariances of the solutions and numerically evaluate the time development for a certain class of curves as in [Ma1, Ma5]; the evaluation is, sometimes, much easier than the approach by means of the theta functions.
@@
In this paper, we leave aside the reality conditions for now and present a generalization of the algebraic differential relation (1.1) of genera two and three, $g=2,3$, in terms of the hyperelliptic sigma functions associated with the hyperelliptic solutions of the NLS equation. To obtain this, we express the Baker function in terms of the sigma function reported in [BEL1], though our Baker function differs from that in [BEL1]. Using the Baker equation, we demonstrate this generalization. We note that the Baker equation, or the generalized Lamé equation, of the Baker function contains the potential $\wp_{gg}$, while the addition theorem of the sigma functions gives the meromorphic functions of the Jacobian $J_{X}$, which does not coincide with $\wp_{gg}$ itself except for $g=1$. Thus, this discrepancy prevents us from obtaining a generalization of (1.1). Even for genera two and three, we find a generalization of the algebraic differential relation (1.1) in this paper. In other words, the purpose of this paper is to show how to overcome the discrepancy so that we can generalize (1.1) associated with the hyperelliptic solution of the NLS equation.
@@
We note that in the NLS equation, we have crucial problems related to the reality conditions depending on the “nonfocusing” and “focusing” NLS equations. These correspond to the sign of the nonlinear terms in the NLS equation [P] and to the choice of $(\alpha,u)$ which belongs to ${\mathfrak{i}}{\mathbb{R}}\times{\mathbb{R}}$ or ×${\mathbb{R}}\times{\mathfrak{i}}{\mathbb{R}}$. However, we will not address these problems at this stage, because we plan to consider them in the next step after considering the algebraic properties of the generalized algebraic differential relation.
@@
The remainder of this paper is organized as follows. Section 2 is devoted to the genus one case or the derivation of (1.1). Section 3 presents a preliminary of the hyperelliptic sigma function. Section 4 provides the addition formula related to the NLS equation. In Section 5, we demonstrate the Baker functions in terms of hyperelliptic sigma functions. In Section 6, we generalize (1.1) for the genera two and three cases using the Baker function and the addition formulae in Theorems 6.2 and 6.3.
@@ 2. Genus one case
(2.1) $E:\ \frac{1}{4}{\widehat{y}}^{2}=y^{2}=x^{3}+\lambda_{2}x^{2}+\lambda_{1}x+\lambda_{0}=(x-e_{1})(x-e_{2})(x-e_{3}),$
where the $e$’s are complex numbers, $\lambda_{2}=-(e_{1}+e_{2}+e_{3})$ and ${\widehat{y}}=2y$. In Weierstrass’s convention, $\lambda_{2}=0$, but we leave it nonzero.
@@
with $x(u)=\wp(u),~{}{\widehat{y}}(u)=\wp^{\prime}(u)$ and $\infty$ is the infinite point of $E$. The Jacobian is given by $J_{E}={\mathbb{C}}/(2{\mathbb{Z}}\omega^{\prime}+2{\mathbb{Z}}\omega^{\prime\prime})$ using the double period $(2\omega^{\prime},2\omega^{\prime\prime})$ for $\omega^{\prime}=\displaystyle{\int_{\infty}^{(e_{1},0)}\frac{dx}{2y}}$ and $\omega^{\prime\prime}=\displaystyle{\int_{\infty}^{(e_{3},0)}\frac{dx}{2y}}$.
@@
Let $(\wp(u),\wp^{\prime}(u))=(x_{1},2y_{1})$ and $(\wp(v),\wp^{\prime}(v))=(x_{2},2y_{2})$. Using ${\frac{d}{du}=2y_{1}\frac{d}{dx_{1}}}$ and ${\frac{d}{dv}=2y_{2}\frac{d}{dx_{2}}}$, we find that (2.4) leads to the following relations:
@@
We also have ${\frac{d\psi(u;v)}{du}}=(\zeta(u+v)-\zeta(u)-\zeta(v))\psi(u;v)$, or
@@
Further, using the Miura relation, $2x+\lambda_{2}+{\widetilde{x}}=\mu^{2}+\mu_{u}$ for ${\mu:=\frac{d}{du}\log\psi(u;v)=\frac{y-{\widetilde{y}}}{x-{\widetilde{x}}}}$, we obtain the static MKdV equation,
@@
$\Psi(u;v,t)$ ::=:= ${\mathrm{e}}^{\alpha u+{\mathfrak{i}}\beta t+c}\psi(u;v)=\frac{{\mathrm{e}}^{\alpha u+{\mathfrak{i}}\beta t+c-\zeta(v)u}\sigma(u+v)}{\sigma(u)\sigma(v)},$
$\Psi^{*}(u;v,t)$ ::=:= ${\mathrm{e}}^{-\alpha u-{\mathfrak{i}}\beta t-c}\psi(u;-v)=-\frac{{\mathrm{e}}^{-\alpha u-{\mathfrak{i}}\beta t-c+\zeta(v)u}\sigma(u-v)}{\sigma(u)\sigma(v)},$
for $\alpha,\beta,c\in{\mathbb{C}}$. Here we use $\sigma(-v)=-\sigma(v)$ and $\zeta(-v)=-\zeta(v)$. We have
$\left[-{\mathfrak{i}}\frac{\partial}{\partial t}+\alpha\frac{\partial}{\partial u}\right]\Psi(u)-\frac{1}{2}\frac{\partial^{2}}{\partial u^{2}}\Psi(u)+\wp(u)\Psi(u)=\frac{1}{2}(2\beta+\alpha^{2}-\wp(v)-\lambda_{2})\Psi(u).$
Then the addition theorem (2.4), $\Psi^{*}\Psi=\wp(u)-\wp(v)$, leads the relation,
(2.14) $\left[-{\mathfrak{i}}\frac{\partial}{\partial t}+\alpha\frac{\partial}{\partial u}\right]\Psi-\frac{1}{2}\frac{\partial^{2}}{\partial u^{2}}\Psi+(\Psi^{*}\Psi)\Psi=\frac{1}{2}(2\beta+\alpha^{2}-3\wp(v)-\lambda_{2})\Psi,$
which is the same as (1.1) when $\lambda_{2}=0$.
@@ 3. Hyperelliptic curve $X$ and sigma functions
In this Section, we provide a background information on the hyperelliptic $\theta$ functions and the $\sigma$ functions as a generalization of the Weierstrass elliptic $\sigma$ function.
@@ 3.1. Geometrical setting of hyperelliptic curves
together with a smooth point $\infty$ at infinity. Here $\lambda$’s are complex numbers. The affine ring related to $X$ is denoted as $R_{X}:={\mathbb{C}}[x,y]/(y^{2}-f(x))$. We fix the basis of the holomorphic one-form
@@
$\omega^{\prime}=\frac{1}{2}\left[\oint_{\alpha_{j}}{\nu^{\mathrm{I}}_{i}}\right],\quad\omega^{\prime\prime}=\frac{1}{2}\left[\oint_{\beta_{j}}{\nu^{\mathrm{I}}_{i}}\right].$
@@
${\widetilde{w}}:S^{k}{\widetilde{X}}\to{\mathbb{C}},\quad{\widetilde{w}}(\gamma_{1},\ldots,\gamma_{k})=\sum_{i=1}^{k}\int_{\gamma_{i}}\left[\begin{array}[]{c}{\nu^{\mathrm{I}}_{1}}\\ \vdots\\ {\nu^{\mathrm{I}}_{g}}\end{array}\right],$
$w:S^{k}X\to J_{X},\quad w((x_{1},y_{1}),\cdots,(x_{k},y_{k}))=\sum_{i=1}^{k}\int_{\infty}^{(x_{i},y_{i})}\left[\begin{array}[]{c}{\nu^{\mathrm{I}}_{1}}\\ \vdots\\ {\nu^{\mathrm{I}}_{g}}\end{array}\right]\hbox{\rm mod}\ {\Lambda}.$
The image of $w$ is denoted by $W_{X}^{k}=w(S^{k}X)$, i.e., $J_{X}=W_{X}^{g}$. Furthermore we introduce an injection $\iota_{X}:X\to{\widetilde{X}}$ and will fix it. We find that $w=\kappa_{J}\circ{\widetilde{w}}\circ\iota_{X}$.
@@
    ${\mathcal{X}}_{g}:={\small{\left(\begin{array}[]{cccc}\chi_{1,0}&\chi_{1,1}&\cdots&\chi_{1,g-1}\\ \chi_{2,0}&\chi_{2,1}&\cdots&\chi_{2,g-1}\\ \vdots&\vdots&\ddots&\vdots\\ \chi_{g,0}&\chi_{g,1}&\cdots&\chi_{g,g-1}\end{array}\right)}},\quad{\mathcal{Y}}_{g}:={\small{\left(\begin{array}[]{cccc}y_{1}&&&\\ &y_{2}&&\\ &&\ddots&\\ &&&y_{g}\end{array}\right)}}=\frac{1}{2}{\mathcal{H}}_{g},$
    ${\mathcal{F}}_{g}^{\prime}:={\small{\left(\begin{array}[]{cccc}F^{\prime}(x_{1})&&&\\ &F^{\prime}(x_{2})&&\\ &&\ddots&\\ &&&F^{\prime}(x_{g})\end{array}\right)}},{\mathcal{V}}_{g}:={\small{\left(\begin{array}[]{cccc}1&1&\cdots&1\\ x_{1}&x_{2}&\cdots&x_{g}\\ x_{1}^{2}&x_{2}^{2}&\cdots&x_{g}^{2}\\ \cdot&\cdot&&\cdot\\ x_{1}^{g-1}&x_{2}^{g-1}&\cdots&x_{g}^{g-1}\end{array}\right)}},$
    ${\mathcal{E}}_{g}:={\small{\left(\begin{array}[]{ccccc}1&&&&\\ \varepsilon_{g-1}&1&&\mbox{\Large 0}&\\ \varepsilon_{g-2}&\varepsilon_{g-1}&1&&\\ \vdots&\vdots&\vdots&\ddots&\\ \varepsilon_{1}&\varepsilon_{2}&\cdots&\varepsilon_{g-1}&1\\ \end{array}\right)}},\quad{\mathcal{K}}_{g}:={\small{\left(\begin{array}[]{cccc}x_{1}^{g-1}&x_{1}^{g-2}&\cdots&1\\ x_{2}^{g-1}&x_{2}^{g-2}&\cdots&1\\ \vdots&\vdots&\ddots&\vdots\\ x_{g}^{g-1}&x_{g}^{g-2}&\cdots&1\end{array}\right)}},$
@@
    The inverse matrix of ${\mathcal{X}}_{g}$ is given as ${\mathcal{X}}_{g}^{-1}={\mathcal{F}}_{g}^{\prime-1}{\mathcal{V}}_{g}$.
@@
    ${\small{\left[\begin{array}[]{c}\partial_{u_{1}}\\ \partial_{u_{2}}\\ \vdots\\ \partial_{u_{g}}\end{array}\right]}}=2{\mathcal{Y}}_{g}{\mathcal{F}}_{g}^{\prime-1}{\mathcal{X}}_{g}{\small{\left[\begin{array}[]{c}\partial_{x_{1}}\\ \partial_{x_{2}}\\ \vdots\\ \partial_{x_{g}}\end{array}\right],}}\quad\frac{\partial x_{i}}{\partial u_{r}}=\frac{2y_{i}}{F^{\prime}(x_{i})}\chi_{i,r-1},\quad\frac{\partial y_{i}}{\partial u_{r}}=\frac{f^{\prime}(x_{i})}{F^{\prime}(x_{i})}\chi_{i,r-1},$